OpenStudy (anonymous):

pls explain this, because i don't get this either. If f(x) = 2x + 6, find f(7). (A)14 (B)15 (C)20 (D)6

OpenStudy (hoblos):

replace 7 in the place of x f(7)= 2(7) +6 = 14+6 =20

OpenStudy (hoblos):

get it ?

OpenStudy (anonymous):

yeah i get man thanks

OpenStudy (hoblos):

good :D

OpenStudy (anonymous):

so that mean that f(x) = −4x + 1, find f(−5). is 20 also right?

OpenStudy (hoblos):

you forgot the 1 it is 21